# Understanding and Mitigating the Threat of Vec2Text to Dense Retrieval Systems

### v1

@@
Abstract
The introduction of Vec2Text, a technique for inverting text embeddings, has raised serious privacy concerns within dense retrieval systems utilizing text embeddings, including those provided by OpenAI and Cohere. This threat comes from the ability for a malicious attacker with access to text embeddings to reconstruct the original text.
In this paper, we investigate various aspects of embedding models that could influence the recoverability of text using Vec2Text. Our exploration involves factors such as distance metrics, pooling functions, bottleneck pre-training, training with noise addition, embedding quantization, and embedding dimensions – aspects not previously addressed in the original Vec2Text paper. Through a thorough analysis of these factors, our aim is to gain a deeper understanding of the critical elements impacting the trade-offs between text recoverability and retrieval effectiveness in dense retrieval systems. This analysis provides valuable insights for practitioners involved in designing privacy-aware dense retrieval systems. Additionally, we propose a straightforward fix for embedding transformation that ensures equal ranking effectiveness while mitigating the risk of text recoverability.
Furthermore, we extend the application of Vec2Text to the separate task of corpus poisoning, where, theoretically, Vec2Text presents a more potent threat compared to previous attack methods. Notably, Vec2Text does not require access to the dense retriever’s model parameters and can efficiently generate numerous adversarial passages.
In summary, this study highlights the potential threat posed by Vec2Text to existing dense retrieval systems, while also presenting effective methods to patch and strengthen such systems against such risks.
1 Introduction
Text embeddings are dense vector representations which capture semantic information about the text they encode Muennighoff et al. (2023). Search engines that leverage these embeddings often employ dense retrievers (DRs) Tonellotto (2022); Zhao et al. (2022); Guo et al. (2022); Bruch (2024). These retrievers utilize text embedding models to encode both queries and documents into embeddings. A similarity metric, such as cosine similarity, is then used to estimate relevance. DRs have demonstrated improved retrieval effectiveness compared to traditional exact term-matching search systems, arguably due to the rich semantic information encoded in the embeddings Yates et al. (2021).
However, a recent study conducted by Morris et al. (2023) raises serious privacy concerns regarding DRs. This study explored the issue of inverting textual embeddings: recovering the original text from its embedding. The proposed Vec2Text method iteratively corrects and generates text to reconstruct the original text based on the given input embedding. Vec2Text can accurately recover 92% of short text and reveal sensitive information (such as patient names in clinical notes) with high accuracy.
@@ 1 Introduction
Our in-depth look at Vec2Text further revealed another vulnerability affecting dense retrievers: Vec2Text can be employed to conduct corpus poisoning attacks on dense retrievers Zhong et al. (2023). A corpus poisoning attack involves a malicious actor generating adversarial passages to maximize similarity with a set of training queries. Adversarial passages are then inserted into a retrieval corpus, effectively tricking the ranker into retrieving them for all unseen user queries, thus undermining the user experience of the targeted search system. Vec2Text presents a theoretically more dangerous method for corpus poisoning compared to previous approaches because it does not require access to the embedding model parameters and can easily generate a large number of adversarial passages. However, there has been no study yet to investigate how Vec2Text performs in corpus poisoning attacks. Therefore, in this paper, we present our preliminary results of applying Vec2Text to the corpus poisoning task. Our findings demonstrate that Vec2Text could pose a serious threat to current DR systems.
Our contributions can be summarized in five main aspects:
  • We reproduce Vec2Text, fixing a serious experimental error in the use of the output layer of the GTR-base embeddings. This fix has implications in terms of both retrieval and reconstructibility effectiveness, although the trends observed in the original work do not change.
@@
  • We are the first to adapt Vec2Text to the corpus poisoning task, empirically demonstrating Vec2Text’s capability to efficiently generate adversarial passages without access to embedding model parameters, posing a serious threat to DR systems’ integrity and user experience.
2 The Vec2Text Method
Figure 1: Overview of Vec2Text, taken from Morris et al. (2023) with permission. “Given access to a target embedding $e$ (blue) and query access to an embedding model $\phi$ (blue model), the system aims to iteratively generate (yellow model) hypotheses $\hat{e}$ (pink) to reach the target. Example input is a [passage] taken from a recent Wikipedia article (June 2023). Vec2Text perfectly recovers this text from its embedding after 4 rounds of correction.”
Vec2Text inverts the text encoding process: given an input embedding, generate the text it represents Morris et al. (2023). (A graphical overview of Vec2Text, taken from Morris et al. (2023) with permission, is shown in Figure 1.) The method comprises two training stages. In the initial stage, a hypothesis text generation model is trained, utilizing a conditional transformer generative model that exclusively takes the embedding as the model input. The training objective is to produce the original text. However, Morris et al. found that this simplistic model is insufficient for generating highly accurate original text. Consequently, they consider the text generated by this first stage model as a hypothesis.
Moving on to the second stage, the learning process involves training another transformer generative model. This model aims to generate satisfactory text by refining the initial hypothesis generated in the first stage. The refinement is achieved through an iterative re-embedding and correction training objective. In each iteration step, the model takes the ground-truth embedding, the generated text, and its embedding from the last iteration step as inputs (the 0 step uses the hypothesis text generated from the first stage model). The output target is the original text. This iterative process allows the model to focus on the differences between the generated text and the original text in the embedding space and gradually reduce these differences.
@@ 2 The Vec2Text Method
During inference, beam search is used to guide the generation. At each step of correction, a number of possible corrections are considered, and the top unique continuations are selected based on their distance in embedding space to the ground-truth embedding.
Morris et al. (2023) trained Vec2Text to invert222i.e. to reconstruct the original text that corresponds to the embeddings. GTR-base Ni et al. (2022) and OpenAI’s text-embeddings-ada-002 embedding models. It achieved high reconstructibility333i.e. the capability to reconstruct the original text from the embedding. on in-domain datasets and adapted well to different-length inputs. The method outperformed baselines on various metrics, and running beam search improved the exact match score. In a case study on MIMIC III clinical notes, Vec2Text demonstrated high performance in reconstructing GTR-embedded clinical notes, re-identifying high percentages of first names, last names, and complete names: this raises serious privacy concerns about the embeddings.
@@
3 Reproduction of vec2text
3.1 Experimental Methodology
@@
We use the official Vec2Text codebase [^4] https://github.com/jxmorris12/vec2text provided by the authors and adhere to the training parameters outlined in the original Vec2Text paper. Specifically, we train Vec2Text on documents from the Natural Question (NQ) dataset Kwiatkowski et al. (2019), setting the max sequence length to 32. The Adam optimizer is employed with a learning rate of $2\times 10^{-4}$, incorporating warmup and linear decay. A batch size of 512 is utilized, and all models are trained on a single H100 GPU. The only deviation from the original setting is the number of training epochs. We note that the complete Vec2Text training pipeline is time-consuming. The authors reported in the paper that the full pipeline takes 2 days to train for 100 epochs with four A6000 GPUs. However, in our implementation using the provided code, the full pipeline takes approximately 10 days with a single H100 GPU. Given the extensive number of embedding models we plan to investigate and the limited computational resources at our disposal, we train for 50 epochs. This reduces the training time to around 5 days for each embedding model, making the process more feasible within our constraints. For the Vec2Text inference, we always use beam search with 50 steps and a beam width 4.
@@
In addition to the variation in the number of epochs, we have identified a significant issue in the original experiment involving the training of Vec2Text for reconstructing text from the GTR-base embeddings. Morris et al. inadvertently used the last hidden layer outputs of the GTR-base model as the final text embeddings. However, it’s crucial to note that GTR-base incorporates an additional dense pooling layer to project the last hidden layer outputs, a detail overlooked by the Vec2Text training code. To address this issue, this section includes a re-training of Vec2Text with the correct GTR-base embeddings, and we compare the results with the original setting. This correction ensures a more accurate evaluation of Vec2Text’s performance in inverting GTR-base embeddings.
@@
Finally, to compare how different embedding models balance retrieval effectiveness and privacy preservation, we also measure the retrieval effectiveness of all considered embedding models. Following the original Vec2Text training setting, which involves training Vec2Text on documents from the NQ dataset, we adopt the top-k retrieval accuracy evaluation setting from DPR Karpukhin et al. (2020), which measures the percentages of retrieved passages that contain a span that answers the question (query). This evaluation is commonly employed in the literature for assessing retrieval effectiveness on the NQ dataset Gao and Callan (2022); Wang et al. (2023); Qu et al. (2021); Ma et al. (2021); Xiong et al. (2020). We use the official NQ test queries and Pyserini Lin et al. (2021) IR toolkit for this evaluation.
@@
To evaluate reconstructibility, i.e. the ability to recover the original text from its embedding, we adopt the same measures used by Morris et al. (2023):
@@
  • BLEU score Papineni et al. (2002), a metric capturing n-gram similarities between the true and reconstructed text;
@@
We randomly sample 1,000 passages from a held-out set of NQ training data for this evaluation [^5] From their code, we understood that also Morris et al. performed sampling; while they do not mention how many passages they sampled, the default value in their code is 1,000 and so we adopted that value..
Table 1: Reproduction results of Vec2Text Morris et al. (2023) comparing retrieval effectiveness and text reconstructibility. Bug fixed represents a correction to ensure Vec2Text uses the final dense pooling layer of GTR-base for output. Subscripts denote statistically significantly difference in retrieval effectiveness using a paired, two-tails t-test ($p<0.01$).
Embedding Models Retrieval effectiveness Reconstructibility
top 10 top 20 top 100 top 1000 bleu tf1 exact cos
a. GTR-base (original) 0.634 0.711 0.816 0.893 90.128 0.970 0.758 0.990
b. GTR-base (reproduced Morris et al. (2023)) 0.634 0.711 0.816 0.893 88.956 0.971 0.699 0.996
[…]
3.2 Reproduction Results
Table 1 shows Vec2Text reproduction results. The first row in the table shows the results obtained using the Vec2Text model checkpoint provided by the authors, which was trained with incorrect GTR-base embeddings. Despite the incorrect embeddings, the results indicate that it still performs well for retrieval, yielding reasonable top-k retrieval accuracy. Additionally, Vec2Text demonstrates robust reconstructibility. It is noteworthy that these outcomes are achieved with the incorrect GTR-base embeddings (without the dense pooling layer), suggesting that the last hidden layer of GTR-base already contains sufficient information for Vec2Text to effectively reconstruct the original input text.
@@ 3.2 Reproduction Results
The second row in the table shows the results obtained by our reproduced Vec2Text model with incorrect GTR-base embeddings (i.e., trying to exactly reproduce the original Vec2Text results). The only difference from the first row is the number of epochs; our reproduction uses half the number of epochs (50 epochs) compared to the original checkpoint (100 epochs). Since our goal is to replicate the original Vec2Text results, we also utilize the incorrect GTR-base embeddings, resulting in exactly the same retrieval effectiveness to the original setting. Regarding Vec2Text reconstructibility, our reproduced checkpoint only shows a slight decrease in the exact match score but achieves very close scores on other evaluation metrics. These results suggest that Vec2Text almost converges with half of the epochs than those used by Morris et al., and the only notable improvement with longer training is in exact reconstruction. From here we only consider Vec2Text models trained on 50 epochs, noting their results are likely similar to those obtained if 100 epochs were used.
@@
Finally, in the last row, we present our results using the correct GTR-base embeddings for retrieval and training Vec2Text with 50 epochs [^6] We have made the checkpoint publicly available on the HuggingFace: https://huggingface.co/ielabgroup/vec2text_gtr-base-st_corrector . Our findings emphasize the significance of using the correct embedding for retrieval, as the top-k accuracy exhibits a substantial increase compared to that obtained with incorrect embeddings. On the other hand, we achieve comparable Vec2Text reconstructibility with the original checkpoint, despite utilizing only half the training epochs. These results suggest that employing the correct outputs from the embedding model, rather than the intermediate layer outputs, may contribute to improved Vec2Text reconstruction.
@@
Table 2: Reproduction of the DPR Karpukhin et al. (2020) embedding models with different distance metrics and pooling functions. "cls" indicates the utilization of the CLS token embedding as the representation, while "mean" indicates the use of mean pooling, aggregating all token embeddings to generate the final text embedding. "dot" indicates the dot product distance metric and "cos" indicates the cosine similarity distance metric. There is no statistically significant difference observed in terms of retrieval effectiveness.
@@
a. DPR_cls_dot Dot-product CLS token 0.748 0.800 0.863 0.914 70.735 0.875 0.319 0.985
b. DPR_cls_cos Cosine similarity CLS token 0.748 0.799 0.863 0.914 71.834 0.879 0.345 0.975
c. DPR_mean_cos Cosine similarity Mean pooling 0.745 0.802 0.866 0.916 81.918 0.924 0.480 0.988
4 Understanding what impacts Vec2Text effectiveness

### v2

@@
Abstract.
The emergence of Vec2Text — a method for text embedding inversion — has raised serious privacy concerns for dense retrieval systems which use text embeddings, such as those offered by OpenAI and Cohere. This threat comes from the ability for a malicious attacker with access to embeddings to reconstruct the original text.
In this paper, we investigate various factors related to embedding models that may impact text recoverability via Vec2Text. We explore factors such as distance metrics, pooling functions, bottleneck pre-training, training with noise addition, embedding quantization, and embedding dimensions, which were not considered in the original Vec2Text paper. Through a comprehensive analysis of these factors, our objective is to gain a deeper understanding of the key elements that affect the trade-offs between the text recoverability and retrieval effectiveness of dense retrieval systems, offering insights for practitioners designing privacy-aware dense retrieval systems. We also propose a simple embedding transformation fix that guarantees equal ranking effectiveness while mitigating the recoverability risk.
Overall, this study reveals that Vec2Text could pose a threat to current dense retrieval systems, but there are some effective methods to patch such systems [^1] Code available at https://github.com/ielab/vec2text-dense_retriever-threat.
Vec2text, privacy protection for dense retriever, embedding models.
ccs: Information systems Language models
1. Introduction
Text embeddings are dense vector representations which capture semantic information about the text they encode (Muennighoff et al., 2023). Search engines that leverage these embeddings often employ dense retrievers (DRs) (Tonellotto, 2022; Zhao et al., 2022; Guo et al., 2022; Bruch, 2024). These retrievers utilize text embedding models to encode both queries and documents into embeddings. A similarity metric (e.g., cosine similarity) is then used to estimate relevance. DRs have demonstrated improved retrieval effectiveness compared to traditional exact term-matching search systems, arguably due to the rich semantic information encoded in embeddings (Yates et al., 2021).
However, a recent study conducted by Morris et al. (2023) raises serious privacy concerns regarding DRs. This study explored the issue of inverting textual embeddings: recovering the original text from its embedding. The proposed Vec2Text method iteratively corrects and generates text to reconstruct the original text based on the given input embedding. According to the results presented in the original paper, Vec2Text can accurately recover 92% of short text and reveal sensitive information (such as patient names in clinical notes) with high accuracy.
@@ 1. Introduction
Figure 1. Overview of Vec2Text, taken from Morris et al. (2023) with permission. “Given access to a target embedding $e$ (blue) and query access to an embedding model $\phi$ (blue model), the system aims to iteratively generate (yellow model) hypotheses $\hat{e}$ (pink) to reach the target. Example input is a [passage] taken from a recent Wikipedia article (June 2023). Vec2Text perfectly recovers this text from its embedding after 4 rounds of correction.”
Our contributions can be summarized in four main aspects:
  1. (1)
    We reproduce Vec2Text, correcting an embedding model implementation error and a testing data leakage error in the original experimental setting. This provides a more accurate evaluation of Vec2Text reconstructibility effectiveness and its relation to the embedding model retrieval effectiveness.
  2. (2)
@@
2. The Vec2Text Method
Vec2Text inverts the text encoding process: given an input embedding, generate the text it represents (Morris et al., 2023). (A graphical overview of Vec2Text, taken from Morris et al. (2023) with permission, is shown in Figure 1.) The method comprises two training stages. In the initial stage, a hypothesis text generation model is trained, utilizing a conditional transformer generative model that exclusively takes the embedding as the model input. The training objective is to produce the original text. However, Morris et al. found that this simplistic model is insufficient for generating highly accurate original text. Consequently, they consider the text generated by this first stage model as a hypothesis.
Moving on to the second stage, the learning process involves training another transformer generative model. This model aims to generate satisfactory text by refining the initial hypothesis generated in the first stage. The refinement is achieved through an iterative feedback, re-embedding and correction process. In each iteration step, the model takes the ground-truth embedding, the generated text, and its embedding from the last iteration step as inputs (the 0 step uses the hypothesis text generated from the first stage model). The output target is the original text. This iterative process allows the model to focus on the differences between the generated text and the original text in the embedding space and gradually reduce these differences.
@@ 2. The Vec2Text Method
During inference, a sequence-level beam search is used to guide the generation. At each step of correction, a number of possible corrections are considered, and the top unique continuations are selected based on their distance in embedding space to the ground-truth embedding.
Morris et al. (2023) trained Vec2Text to invert333i.e. to reconstruct the original text that corresponds to the embeddings. GTR-base (Ni et al., 2022) and OpenAI’s text-embeddings-ada-002 embedding models. It achieved high reconstructibility444i.e. the capability to reconstruct the original text from the embedding. on in-domain datasets and adapted well to different-length inputs. The method outperformed baselines on various metrics, and running beam search improved the exact match score. In a case study on MIMIC III clinical notes, Vec2Text demonstrated high performance in reconstructing GTR-embedded clinical notes, re-identifying high percentages of first names, last names, and complete names: this raises serious privacy concerns about the embeddings.
@@
3. Reproduction of vec2text
3.1. Experimental Methodology
Table 1. Reproduction of Vec2Text’s text reconstructibility results on NQ dataset. Bug fixed represents a correction to ensure Vec2Text uses the final dense pooling layer of GTR-base for output.
Model Checkpoints Original NQ valid set Filtered NQ valid set
bleu tf1 exact cos bleu tf1 exact cos
1. GTR-base (paper  (Morris et al., 2023)) 97.3 99.0 92.0 0.99 - - - -
2. GTR-base (author checkpoint) 98.2 99.5 94.4 0.998 97.3 99.2 92.2 1.0
3. GTR-base (reproduced) 97.0 99.2 91.1 1.000 96.1 98.9 89.0 1.0
4. GTR-base (bug fixed) 97.9 99.4 94.0 1.000 96.4 99.0 90.1 1.0
Table 2. Retrieval effectiveness of GTR-base with/without the last dense pooling layer. * denote statistically significantly difference using a paired, two-tails t-test ($p<0.01$).
GTR-base top 10 top 20 top 100 top 1000
Without pooling layer 0.634 0.711 0.816 0.893
With pooling layer 0.718 0.773 0.855 0.913
@@ 3.1. Experimental Methodology
We use the official Vec2Text codebase [^5] https://github.com/jxmorris12/vec2text provided by the authors and adhere to the training parameters outlined in the guidance of the code repository. Specifically, we train Vec2Text on documents from the Natural Question (NQ) dataset (Kwiatkowski et al., 2019), setting the max sequence length to 32. The AdamW optimizer is employed with a learning rate of $1\times 10^{-3}$, incorporating warmup. A batch size of 512 is utilized, and all models are trained on a single H100 GPU for 50 epochs. One major deviation from the original setting is the number of training epochs. We note that the complete Vec2Text training pipeline is time-consuming. The authors reported in the paper that the full pipeline takes 2 days to train for 100 epochs with four A6000 GPUs. However, in our implementation using the provided code, the full pipeline takes approximately 10 days with a single H100 GPU. Given the extensive number of embedding models we plan to investigate and the limited computational resources at our disposal, we train for 50 epochs. This reduces the training time to around 5 days for each embedding model, making the process more feasible within our constraints.
In addition to the variation in the number of epochs, we have identified two significant issues in the original experimental setting.
  1. 1)
@@
    GTR-base model missing dense pooling layer: The first issue involves the training of Vec2Text for reconstructing text from the GTR-base embeddings. Morris et al. inadvertently used the last hidden layer outputs of the GTR-base model as the final text embeddings. However, it’s crucial to note that GTR-base model incorporates an additional dense pooling layer to project the last hidden layer outputs, a detail overlooked by the Vec2Text training code. To address this issue, this section includes a re-training of Vec2Text with the correct GTR-base embeddings, and we compare the results with the original setting.
  2. 2)
    Data leakage in NQ dataset splits: The second issue relates to a serious data leakage problem in the author-provided Vec2Text NQ training and validation dataset splits. We found that $50.7\%$ of validation data points also appear in the training dataset. Although the authors evaluated Vec2Text also on out-domain datasets and demonstrated its reconstruction effectiveness, the reported effectiveness on the NQ dataset (in-domain evaluation) reported in the original paper was likely overestimated. Hence, in our experiments, we re-evaluate Vec2Text for the NQ datasets with a corrected, non-overlapping validation set.
These two corrections ensure a more accurate evaluation of Vec2Text’s performance in inverting GTR-base embeddings.
@@
Finally, to compare how different embedding models balance retrieval effectiveness and privacy preservation (i.e., not being reconstructed by Vec2Text), we also measure the retrieval effectiveness of all considered embedding models. Following the original Vec2Text training setting, which involves training Vec2Text on documents from the NQ dataset, we adopt the top-k retrieval accuracy evaluation setting from DPR (Karpukhin et al., 2020), which measures the percentages of retrieved passages that contain a span that answers the question (query). This evaluation is commonly employed in the literature for assessing retrieval effectiveness on the NQ dataset (Gao and Callan, 2022; Wang et al., 2023; Qu et al., 2021; Ma et al., 2021; Xiong et al., 2020). We use the official NQ test queries and Pyserini (Lin et al., 2021) IR toolkit for this evaluation.
@@
To evaluate reconstructibility (i.e., the ability to recover the original text from its embedding) we adopt the same measures used by Morris et al. (2023):
@@
  • BLEU score (Papineni et al., 2002), a metric capturing n-gram similarities between the true and reconstructed text;
@@
We follow the default evaluation setting in the original codebase, which requires the random sampling of 1,000 passages from the full NQ validation set. We note that the full size of the author-provided NQ validation set has 849,508 data points, of which 50.7% also appear in the training set. To resolve this data leakage problem, we further filtered out these overlapping data points to create our filtered NQ validation set, which now consists of 419,851 data points.
For the Vec2Text generation inference configuration, we use sequence-level beam search with 50 steps and a beam width of 8. The whole evaluation for a single experiment took around 4 hours on one Nvidia H100 GPU.
3.2. Reproduction Results
Table 1 shows Vec2Text reproduction results on Morris et al.’s NQ validation set and our filtered NQ validation set. For reference, in the first row of the table we list scores as appear in the original paper. We note that these scores are obtained on Morris et al.’s validation set which is affected by the data leakage.
We first consider evaluating the author-provided Vec2Text model checkpoint, which is supposed to be the same as the checkpoint evaluated and reported in the original paper. This experiment aims to verify the correctness of our evaluation setting and understand the impact of the data leakage issue. We report our results in the second row of Table 1. Compared to the scores reported by Morris et al., we obtained higher scores on the original validation split, likely due to the randomness of the sampled subset. The evaluation scores on our filtered validation set are only slightly lower, suggesting that the author-provided checkpoint did not overfit the training data and generalizes well to the non-overlapping data points. Thus, although there is a considerable amount of data leakage, the negative impact on the evaluation validity appears to be negligible.
@@ 3.2. Reproduction Results
In the third row, we report the results obtained from our reproduced Vec2Text model with incorrect GTR-base embeddings (i.e., trying to exactly reproduce the original Vec2Text results). We note that the major training difference from the original checkpoint is the number of training epochs; our reproduction uses half the number of epochs (50 epochs) compared to the original checkpoint (100 epochs). Despite this difference, our reproduced checkpoint only shows a slight decrease in the BLEU and exact match scores but achieves very close scores on other evaluation metrics on the original validation set, and, like the original checkpoint, only slightly lower on the filtered validation set. These results suggest that Vec2Text almost converges with half of the epochs used by Morris et al., and the only notable improvement with longer training is in exact reconstruction. From here, we only consider Vec2Text models trained for 50 epochs, noting that their results are likely similar to those obtained if 100 epochs were used.
@@
Finally, in the last row, we present the results using the correct GTR-base embeddings for training Vec2Text with 50 epochs. We achieve comparable Vec2Text reconstructibility with the original checkpoint, despite utilizing only half the training epochs. These results suggest that employing the correct outputs from the embedding model, rather than the intermediate layer outputs, may contribute to improved Vec2Text reconstruction. On the other hand, in Table 2 we report the retrieval effectiveness of the GTR-base embedding model with (correct embeddings) and without (incorrect embeddings) the last dense pooling layer. Our results emphasize the significance of using the correct embedding for retrieval, as the top-k accuracy exhibits a substantial increase compared to that obtained with incorrect embeddings.
To conclude this section, our reproduction results demonstrate that using the correct GTR-base embeddings and a reduced number of training epochs can achieve high reconstructibility and retrieval effectiveness, highlighting the robustness and efficiency of the Vec2Text model under different configurations. Additionally, the slightly lower evaluation scores on our filtered validation set suggest that the Morris et al.’s checkpoint generalizes well and that the impact of data leakage is minimal, despite the amount of leakage itself being considerable. For the rest of the experiments in this study, we use our filtered validation set to evaluate the reconstructibility of Vec2Text models.
@@
Table 3. Reproduction of the DPR (Karpukhin et al., 2020) embedding models with different distance metrics and pooling functions. "cls" indicates the utilization of the CLS token embedding as the representation, while "mean" indicates the use of mean pooling, aggregating all token embeddings to generate the final text embedding. "dot" indicates the dot product distance metric and "cos" indicates the cosine similarity distance metric. There is no statistically significant difference observed in terms of retrieval effectiveness.
@@
a. DPR_cls_dot Dot-product CLS token 0.748 0.800 0.863 0.914 79.2 90.9 43.0 0.996
b. DPR_cls_cos Cosine similarity CLS token 0.748 0.799 0.863 0.914 82.3 92.1 50.0 0.994
c. DPR_mean_cos Cosine similarity Mean pooling 0.745 0.802 0.866 0.916 89.4 95.2 57.9 0.996
4. Understanding what impacts Vec2Text effectiveness